# Point Prompting: Counterfactual Tracking with Video Diffusion Models

### v1

@@ Abstract
Trackers and video generators solve closely related problems: the former analyze motion, while the latter synthesize it. We show that this connection enables pretrained video diffusion models to perform zero-shot point tracking by simply prompting them to visually mark points as they move over time. We place a distinctively colored marker at the query point, then regenerate the rest of the video from an intermediate noise level. This propagates the marker across frames, tracing the point’s trajectory. To ensure that the marker remains visible in this counterfactual generation, despite such markers being unlikely in natural videos, we use the unedited initial frame as a negative prompt. Through experiments with multiple image-conditioned video diffusion models, we find that these “emergent” tracks outperform those of prior zero-shot methods and persist through occlusions, often obtaining performance that is competitive with specialized self-supervised models.
@@ 3.3 Extensions
Since our tracker relies on detecting a particular color, we rebalance the video’s colors such that the marker’s color does not appear within it. We do this by reducing the saturation of the marker’s color. For example, when tracking a red marker, we reduce the saturation of red regions, effectively suppressing natural red hues while preserving overall image quality (details provided in Appendix C.1). We find that this reduces mistakes during occlusion, since the marker is not present and thus false detections are more common.
@@ 7 Acknowledgements
Appendix A Limitations
Appendix B Quantitative Results on TAP-Vid
@@ B.2 V-Bench scores
Appendix C Implementation Details
C.1 Video Preprocessing
@@ C.2 Point Propagation
C.3 Tracker
C.3.1 Marker Detection
@@
C.3.2 Local Search and Occlusion Handling

### v2

@@ Abstract
Trackers and video generators solve closely related problems: the former analyze motion, while the latter synthesize it. We show that this connection enables pretrained video diffusion models to perform zero-shot point tracking by simply prompting them to visually mark points as they move over time. We place a distinctively colored marker at the query point, then regenerate the rest of the video from an intermediate noise level. This propagates the marker across frames, tracing the point’s trajectory. To ensure that the marker remains visible in this counterfactual generation, despite such markers being unlikely in natural videos, we use the unedited initial frame as a negative prompt. Through experiments with multiple image-conditioned video diffusion models, we find that these “emergent” tracks outperform those of prior zero-shot methods and persist through occlusions, often obtaining performance that is competitive with specialized self-supervised models. Finally, we show that trajectories produced by pretrained generators can be distilled into a fast tracker with similar performance, serving as effective supervision for a tracking model.
@@ 1 Introduction
  • Trajectories produced by pretrained video generators can be distilled into a fast tracker with comparable performance.
@@ 3.3 Extensions
Since our tracker relies on detecting a particular color, we rebalance the video’s colors such that the marker’s color does not appear within it. We do this by reducing the saturation of the marker’s color. For example, when tracking a red marker, we reduce the saturation of red regions, effectively suppressing natural red hues while preserving overall image quality (details provided in Appendix B.1). We find that this reduces mistakes during occlusion, since the marker is not present and thus false detections are more common.
@@ 5 Results
Method TAP-Vid DAVIS
AJ $\uparrow$ $\delta^{x}_{\textrm{avg}} \uparrow$ OA $\uparrow$
Wan2.1-1.3B (teacher) 38.78 54.75 85.00
Cotracker (distilled) 37.17 53.12 84.24
Table 5: Distilling the generator. We distill our method to CoTracker which achieves performance close to the teacher Wan2.1 and runs orders of magnitude faster.
Distilling the Generator into a Tracker. While our approach shows strong zero-shot tracking performance, it requires a separate video generation for each query point, which limits efficiency. To obtain a fast tracker, inspired by this (Stojanov et al., 2025), we distill our generation-based method into an efficient tracking architecture. Specifically, we collect pseudo-label trajectories by running our marker propagation method on 1,000 unlabeled videos from the TAP-Vid Kinetics dataset. These extracted trajectories serve as supervision to train CoTracker (Karaev et al., 2024d) from scratch.
Table 5 compares the teacher model used to generate trajectories with the distilled CoTracker model. Despite being trained solely on pseudo-labels produced by our zero-shot method, the distilled tracker achieves performance very close to the teacher while running orders of magnitude faster at inference. This indicates that the temporal reasoning capabilities of large video diffusion models can be transferred into a lightweight tracking network, effectively bridging generative zero-shot tracking and efficient feed-forward inference.
@@ 8 Acknowledgements
Appendix A Quantitative Results on TAP-Vid
@@ A.2 V-Bench scores
Appendix B Implementation Details
B.1 Video Preprocessing
@@ B.2 Point Propagation
B.3 Tracker
B.3.1 Marker Detection
@@
B.3.2 Local Search and Occlusion Handling